OpenStudy (0487308):

Frankie can build a speaker box in 120 minutes. Tisha can build a speaker box 40 minutes faster. If they work together, how many minutes does it take them to build the speaker box?

OpenStudy (0487308):

@zepdrix

OpenStudy (kendricklamar2014):

Frank can build a speaker in 120 minutes In 1 minute , Frank can build \[\frac{ 1 }{ 120 }\]th of a speaker Tisha can build a speaker in 80 minutes In 1 minute , Tisha can build \[\frac{ 1 }{ 80 }\]th of a speaker Together, in one minute they both can build \[\frac{ 1 }{ 120 } + \frac{ 1 }{ 80 } = \frac{ 1 }{ 48 }\] of a speaker.

OpenStudy (0487308):

Thank you.

OpenStudy (kendricklamar2014):

(1/120+1/80 = (80+120) / 9600 = 200/9600 = 2/96=1/48 ) \[\frac{ 1 }{ 120 } + \frac{ 1 }{ 80 } = \frac{ (80 + 120) }{ 9600 } = \frac{ 200 }{ 9600}= \frac{ 2 }{ 96 } = \frac{ 1 }{ 48 }\]

OpenStudy (0487308):

Micah can fill a box with books in 15 minutes. While Micah fills the box, Sydney takes the books out puts them on a shelf. With Sydney taking books out of the box, it takes 20 minutes for Micah to fill the box. Which of the following can be used to determine the amount of time it takes for Sydney to empty the box if Micah is not adding books? Is the answer to this 1/15 - 1/x = 1/20?

OpenStudy (kendricklamar2014):

Answer: It would take 48 minutes if they work together to build a speaker.

OpenStudy (kendricklamar2014):

One question per post @0487308

OpenStudy (0487308):

Do you want me to ask that question in a different post?

OpenStudy (kendricklamar2014):

Please do